# HEAM : Hashed Embedding Acceleration using Processing-In-Memory

### 21 Feb 2024 v2

@@ 5. Experimental Setup
Simulation Methodology. We modified DRAMsim3 (li2020dramsim3, ), a cycle-accurate memory simulator, to implement and evaluate the performance of the HEAM system. HBM2 x128 and DDR4-3200 x8 were utilized as base memory systems for the HEAM system and the baselines. Table 2 summarizes the memory system configuration utilized in our experiment. Within a single DDR4-3200 x8 device, 2 ranks were employed. To utilize the heterogeneous memory systems, HEAM makes use of a flat addressable method, which has been continuously adopted in prior works on the same system configuration (kal2021space, ), (sim2014transparent, ), (chou2014cameo, ), (chou2015bear, ). We compare HEAM performance against previous NMP architectures, TensorDIMM (kwon2019tensordimm, ), RecNMP (ke2020recnmp, ), TRiM (park2021trim, ) and SPACE (kal2021space, ). The traces of embedding operation were extracted from running the DLRM model, which is fed into the simulation framework.
In order to meet fair comparison, we treat all the embeddings within the R table as hot vectors so that temporal locality utilization techniques proposed in prior works could be maximally exploited. For instance, we mark every R table embedding with localityBit when implementing RecNMP. For TRiM implementation, R table embeddings have the highest priority when participating in hot entry profiling. In this case, vectors counted from the front of q vectors that are ordered with highest locality takes participation in hot entry profiling after R table. To fairly measure the performance of SPACE, reduction locality technique was employed, storing partial sums of hot vectors in the available space on HBM2.
Area and Power Consumption. To obtain the area and the power of PIM units, bd-PIM and bg-PIM were synthesized using Synopsys Design Compiler with 45nm CMOS technology operating at 300MHz clock frequency. For the case of bg-PIM, the corresponding area was scaled to a 20nm DRAM processing. Area and the power of LUT used inside bg-PIM were estimated using CACTI (muralimanohar2009cacti, ).
Recommendation System Datasets. Obtaining a sufficiently large dataset is crucial for our experiment as HBM2 capacity needs to be fully occupied even after compositional embedding is applied regarding the assumption of the HEAM system. However, since datasets meeting such constraints is not publicly available, we utilize criteo datasets (criteo_kaggle_ad_dataset, ),(reddi2020mlperf, ) by setting its vector size large enough, 128B to 512B. Note that criteo-terabyte dataset (reddi2020mlperf, ) capacity is near 10 GB when the vector length is 64B.
Table 2. Memory System configurations
@@
6. Evaluation Result
@@
We evaluate the impact of HEAM’s design choices as the first step to analyze the effectiveness of each choice. We then compare the performance of HEAM with state-of-the-art NMP designs, which target the embedding operation. Additionally, hardware overhead introduced to the overall system for HEAM architecture support is analyzed. We assess the limitations of compositional embedding to explore the boundary in terms of compositional embedding could be applied to our system in terms of preserving model quality.
6.1. Performance Impact of Design Optimizations
@@
Following TRiM (park2021trim, ), we demonstrate the impact of each design optimization choice on the speedup performance. Figure 9 displays the performance improvements when design optimizations are gradually incorporated in HEAM system. The baseline is the execution speed of HBM2 in conjunction with bd-PIM. The first scenario is HEAM-BG, where bg-PIM without LUT is located in each bank group. The following scheme is CA compression, when command address compression is applied to the HEAM-BG scenario. Batching corresponds to when the batching technique is applied to the HEAM system. Finally, LUT scheme adds LUT to Batching scenario, reducing direct access to the bank of R table. Applying all the optimization techniques achieves 2.8×2.8$\times$2.8 × higher result than the baseline when the vector size is 512B. HEAM-BG enhances the performance by 10% compared to the baseline. CA compression enhances 14% of the performance as the CA bandwidth overhead is mitigated by both utilizing the CA pin and DQ pin of HBM2. The batching scheme gains 86% of additional speedup by reducing load imbalance issues and satisfying the high demand of pooling size of bg-PIM utilized inside HBM2. The root of the performance benefit of Batching lies in pulling up the potential of bank group level parallelism. Finally, 19% more speedup is acquired in the LUT scenario, as it reduces the total DRAM access count by effectively utilizing the high temporal locality of the R table. Discussed highlights of each optimization technique increase as the vector length grows, leading to more frequent ACTs within each bank group, creating an environment that fully allows bg-PIM to capitalize its parallelism.
The primary reason for the relatively low speedup of HEAM-BG, even though it utilizes bg-PIM, is that the total embedding lookups in a single batch are small in the criterion dataset, resulting in low utilization of bg-PIM parallelism. The fact that baseline HBM2 is capable of embedding lookups of criteo dataset due to its abundant channels further diminishes effect of HEAM-BG in this case. Coupled with the operation cycles of PIM and cycles required to deliver PIM instructions, this under-utilization issue especially worsens for small vector length. However, when batching is applied, total embedding lookups per operation increase, resulting in higher throughput than baseline.
Figure 9. Performance Impact of Design Optimizations
6.2. Performance Evaluation with Previous NMP Designs
@@
Speedup Comparison. To demonstrate the effectiveness of our proposal, we compare the overall performance of HEAM to the speedup in embedding operation of the previous works (kwon2019tensordimm, ), (ke2020recnmp, ), (park2021trim, ), (kal2021space, ). In our experiment, a total of two DIMMs were utilized for previous DIMM-based NMP architectures.
@@
Through the use of HBM, 2-level PIM structure, and offloading techniques, HEAM achieved the best speedup across all the prior works. Figure 10 shows the relative performance of configurations normalized to TensorDIMM, varying vector lengths from 128B to 512B. Compared to the prior works, HEAM achieves 6.3×6.3$\times$6.3 ×, 4.6×4.6$\times$4.6 ×, 2.3×2.3$\times$2.3 ×, 1.8×1.8$\times$1.8 × speedup compared to TensorDIMM, RecNMP, TRiM-G, SPACE when vector length is 512B. Compared to SPACE, which also utilizes a heterogeneous memory system, HEAM’s superior achievement in speedup comes from the utilization of bg-PIM followed by design choices of batching and LUT. The performance enhancement of HEAM increases as vector length gets larger, as internal bandwidth becomes more utilized due to more frequent ACTs occurring inside the memory. A larger pooling size and larger vector length, a normal trend of recommendation systems in production, would result in even more enhancement.
Figure 10. Performance evaluation on various NMP systems
@@
Figure 11. Relative energy consumption in NMP models
6.3. Design Overhead
@@
6.4. Assessment of Compositional Embedding Shortcomings
@@
Efficiency of Increasing Hash Collision. To achieve the necessary bandwidth ratio between HBM and DIMM, reducing the total number of hot vectors should coincide with an increase in hash collisions, especially since the total hot vectors might surpass the capacity of a single HBM stack in production scenarios. Ideally, hot vectors would be grouped closely, allowing a quotient hash function to map consecutive hot vectors to the same Q vector. This would inversely reduce the total number of hot vectors relative to the increase in hash collisions. However, this ideal situation is rare. In the worst-case scenario, the reduction in total hot vectors is minimal, as cold vectors typically outnumber and surround each hot Q vector, leading to only a marginal decrease.
@@
To verify the reduction efficiency, we measured the changes in the number of the hot vectors while modifying hash collision value on public datasets available online (criteo_kaggle_ad_dataset, ), (movielens, ). In this experiment, we defined hot vectors as the vectors needed to achieve 80% of the cumulative request. The result is shown in Figure 12 (a). Although the result indicates the hot vectors are not clustered ideally, increasing hash collision still decreases hot vectors, which is enough to meet the bandwidth ratio demand between HBM and DIMM.
@@
Figure 12. Compositional embedding shortcoming analysis. (a) Change in the number of hot vectors based on the hash collision value across criteo-kaggle, movielens-25M, amazon electronics. (b) Accuracy difference between the original model and the model trained with compositional embedding by hash collision 4.
@@
Decrease in model quality drop with more embedding tables. As more embedding tables are used in the recommendation system, the model quality increases (zhao2022understanding, ). At the production level, it is known that there are 10s of embedding tables. As the accuracy drop of compositional embedding is due to the mixture of distinctive embeddings (lyu2022optembed, ), applying more embedding tables might resolve the accuracy drop caused when applying the algorithm. To see the effect of embedding tables on the accuracy, we conducted experiments in criteo-kaggle-dataset. Since some tables are used more often than others, we picked randomly the tables to be excluded, averaging the result of 3 experiments. The overall result is shown in Figure 12 (b). By excluding tables from 10 to 0, we found that the gap of accuracy drop decreases.
7. Conclusions
@@
We introduce HEAM, a 3-level PIM architecture that comprises DIMM, HBM featuring bd-PIM and bg-PIM with LUT. Our goal is to accelerate compositional embedding. We have examined the distinct traits of compositional embedding and leveraged our findings to boost overall throughput. Our system tackles two key challenges in DLRM: model size expansion and memory-bound operations. In addition, HEAM effectively resolves the issue of double memory access that arises when applying compositional embedding, resulting in 6.3×6.3$\times$6.3 × performance improvement and 58.9% energy savings compared to the previous NMP architectures tailored for recommendation systems.

### 14 Mar 2024 v3

@@ V Experimental Setup
Simulation Methodology. We modified DRAMsim3 [27], a cycle-accurate memory simulator, to implement and evaluate the performance of the HEAM system and the prior works. HBM2 x128 and DDR4-3200 x8 were utilized as base memory systems for the HEAM system and the baselines. Table II summarizes the memory system configuration utilized in our experiment. Within a single DDR4-3200 x8 device, 2 ranks were employed. To utilize the heterogeneous memory systems, HEAM makes use of a flat addressable method, which has been continuously adopted in prior works on the same system configuration [9], [28], [29], [30]. We compare HEAM performance against previous NMP architectures, TensorDIMM [6], RecNMP [7], TRiM [8] and SPACE [9]. The traces of embedding operation were extracted from running the DLRM model, which is fed into the simulation framework.
TABLE II: Memory System configurations
@@
In order to meet fair comparison, we treat all the embeddings within the R table as hot vectors so that temporal locality utilization techniques proposed in prior works could be maximally exploited. For instance, R table embeddings have the highest priority when participating in hot entry profiling when implementing TRiM. In this case, vectors counted from the front of q vectors that are ordered with highest locality takes participation in hot entry profiling after the R table. To fairly measure the performance of SPACE, reduction locality technique was employed, storing partial sums of hot vectors in the available space on HBM2. For RecNMP, we linearly scaled the speedup of RankCache, following TRiM.
Area and Power Consumption. To obtain the area overhead and the power consumption of PIM units, bd-PIM and bg-PIM were synthesized using Synopsys Design Compiler with 45nm CMOS technology operating at 300MHz clock frequency. For the case of bg-PIM, the corresponding area was scaled to a 20nm DRAM processing. Area and the power of LUT were estimated using CACTI [31].
Recommendation System Datasets. Obtaining a sufficiently large dataset is crucial for our experiment as HBM2 capacity needs to be fully occupied even after compositional embedding is applied regarding the assumption of the HEAM system. However, since datasets meeting such constraints is not publicly available, we generate synthetic traces from criteo datasets [25], [26] by adjusting their vector sizes between 128B and 512B. Total embedding lookup per operation was set to 78. Note that criteo-terabyte dataset [26] capacity is near 10 GB when the vector length is 64B.
VI Evaluation Result
@@
We evaluate the impact of HEAM’s design choices as the first step to analyze the effectiveness of each choice. We then compare the performance of HEAM with state-of-the-art NMP designs, which target the embedding operation. Additionally, area overhead and power consumption introduced to the overall system for HEAM architecture support are analyzed. Furthermore, we assess the limitations of compositional embedding to explore the boundary in terms of compositional embedding could be applied to our system in terms of preserving model quality.
VI-A Performance Impact of Design Optimizations
@@
Following TRiM [8], we demonstrate the impact of each design optimization choice on the speedup performance. Figure 9 displays the performance improvements when design optimizations are gradually incorporated in HEAM system. The baseline is the execution speed of HBM2 in conjunction with bd-PIM, with command address compression technique being applied. The first scenario is HEAM-BG, where bg-PIM without LUT is located in each bank group. Batching corresponds to when the batching technique is applied to the HEAM system. Finally, LUT scheme adds LUT to Batching scenario, reducing direct access to the bank of R table. Applying all the optimization techniques achieves $\times2$ higher result than the baseline when the vector size is 512B. HEAM-BG enhances the performance by 34% compared to the baseline. The batching scheme gains 43% of additional speedup by reducing load imbalance issues and satisfying the high demand of pooling size that of the bg-PIM. The root of the performance benefit of Batching lies in pulling up the underused potential of bank-group-level parallelism. Finally, 14% more speedup is acquired in the LUT scenario, as it reduces the total DRAM access count by effectively utilizing the high temporal locality of the R table. Discussed highlights of each optimization technique increase as the vector length grows, leading to more frequent ACTs within each bank group, creating an environment that fully allows bg-PIM to capitalize its parallelism.
The relatively low speedup with a small vector size primarily results in the underuse of bg-PIM’s parallel processing capabilities, stemming from the low ACT frequency. This impact is further amplified on HEAM-BG’s speedup, as HBM can already conduct parallel embedding lookups effectively for small vector sizes since it possesses abundant channels. However, when batching is applied, total embedding lookups per PIM operation increase, resulting in higher throughput than baseline.
Figure 9: Performance Impact of Design Optimizations
VI-B Performance Evaluation with Previous NMP Designs
@@
Speedup Comparison. To demonstrate the effectiveness of our proposal, we compare the overall performance of HEAM to the speedup in embedding operation of the previous works [6], [7], [8], [9]. In our experiment, a total of two DIMMs were utilized for previous DIMM-based NMP architectures.
@@
Through the use of a heterogeneous memory system, 2-level PIM structure, and offloading techniques, HEAM achieved the best speedup across all the prior works. Figure 10 shows the relative performance of configurations normalized to TensorDIMM, varying vector lengths from 128B to 512B. Compared to the prior works, HEAM achieves 6.2×6.2$\times$6.2 ×, 4.77×4.$7$, $\times3$, 2.13×2.13$\times$2.13 × speedup compared to TensorDIMM, RecNMP, TRiM-G, SPACE when vector length is 512B. Compared to SPACE, which also utilizes a heterogeneous memory system, HEAM’s superior achievement in speedup comes from the utilization of bg-PIM followed by design choices of batching and LUT. However, bank group parallelism’s impact is minimal for short vector lengths due to fewer vector reads per channel, especially when compared to the nearly uniform speedup of the TRiM. This occurs because HBM2 has four times more channels than two DIMMs, leading to underutilized bank group parallelism. With larger vector sizes, HEAM’s performance significantly improves as the internal bandwidth is better utilized, thanks to more frequent ACT commands. Employing a larger pooling size and larger vector length, a typical trend of recommendation systems in production, would result in full utilization of HEAM’s maximal bandwidth.
@@
Figure 10: Performance evaluation on various NMP systems
Figure 11: Relative energy consumption in NMP models
VI-C Design Overhead
@@
VI-D Assessment of Compositional Embedding Shortcomings
@@
Efficiency of Increasing Hash Collision. To achieve the necessary bandwidth ratio between HBM and DIMM, reducing the total number of hot vectors should coincide adequately with increased hash collisions. If not sufficiently reduced, the total hot vectors might surpass the capacity of a single HBM stack in production scenarios. For an ideal case, hot vectors would be grouped closely, allowing a quotient hash function to map consecutive hot vectors to the same Q vector. This would reduce the total number of hot vectors inversely proportional to the increase in hash collisions. However, this ideal situation is rare. In the worst-case scenario, the reduction in total hot vectors is minimal, as cold vectors typically outnumber and surround each hot Q vector, leading to only a marginal decrease.
@@
To verify the reduction efficiency, we measured the changes in the number of the hot vectors while modifying hash collision value on public datasets available online [25], [32], [33]. In this experiment, we defined hot vectors as the vectors needed to achieve 80% of the cumulative request. The result is shown in Figure 12 (a). Although the result indicates the hot vectors are not clustered ideally, increasing hash collision still decreases hot vectors to a certain extent, which is enough to meet the bandwidth ratio demand between HBM and DIMM.
@@
Figure 12: Compositional embedding shortcoming analysis. (a) Change in the number of hot vectors based on the hash collision value across criteo-kaggle, movielens-25M, amazon electronics. (b) Accuracy difference between the original model and the model trained with compositional embedding by hash collision 4.
@@
Decrease in Model Quality Drop with More Embedding Tables. As more embedding tables are used in the recommendation system, the model quality increases [5]. At the production level, it is known that there are 10s of embedding tables. As the accuracy drop of compositional embedding is due to the mixture of distinctive embeddings [22], applying more embedding tables might resolve the accuracy drop caused when applying the algorithm. To see the effect of embedding tables on the accuracy, we conducted experiments in criteo-kaggle-dataset. Since some tables are used more often than others, we picked randomly the tables to be excluded, averaging the result of 3 experiments. The overall result is shown in Figure 12 (b). By excluding tables from 10 tables to 0 tables, we found that the gap of accuracy drop decreases.
VII Conclusion
@@
We introduce HEAM, a three-tiered memory architecture that comprises DIMM, HBM featuring bd-PIM and bg-PIM and LUT inside bg-PIM. Our system tackles two key challenges in deep learning recommendation system: Model size expansion and Memory-bound operations. We examined the unique characteristics of compositional embedding and leveraged our findings to boost overall throughput. HEAM effectively resolves the issue of double memory access that arises when applying compositional embedding, resulting in 6.2×6.2$\times$6.2 × performance improvement and 58.9% energy savings compared to the previous NMP architectures tailored for recommendation systems.